### v1

@@ 3.2 Low-complexity regions in a gorilla genome
To further confirm this observation, we extracted 298.8Mb of regions in the gorilla genome that are $\geq 10$kb in length without any 51-mer exact matches to 472 human genomes (Li, , 2024). 99.7% of them are marked as LC regions by longdust and none of them are alpha satellite. 95.8% of these gorilla-specific regions are distributed within 15Mb from telomeres, broadly in line with Yoo et al., (2025). We also ran TRF on the gorilla genome. It did not finish in 30 hours even with option “-l20” which is supposed to reduce runtime.
@@ 4 Discussions
Implemented in the C programming language, longdust is a fast and lightweight command-line tool for identifying low-complexity regions. It rarely finds LC regions not reported by TRF plus SDUST and can recover more tandem repeats with $\geq 4$ copies of repeat units. Longdust provides basic APIs in C and can also be used as a programming library.
@@
A major practical limitation of longdust is the restricted window size. The genome of Woodhouse’s scrub jays, for example, contains satellite with a 12kb repeat unit (Edwards et al., , 2025). This would be missed by longdust under the default setting. Increasing the window size would make longdust considerably slower. This is partly due to the $O(wL)$ time complexity and partly due to the speedup strategies in longdust that are more effective given $\ell(x)\ll 4^{k}$. It would be ideal to have an algorithm that remains efficient given large windows or, better, does not require specified window sizes.
@@ Acknowledgments
We would like to thank Qian Qin for evaluating the effect of low-complexity regions in structural variant calling.

### v2

@@ 3.4 Low-complexity regions in a gorilla genome
To further confirm this observation, we extracted 298.8Mb of regions in the gorilla genome that are $\geq 10$kb in length without any 51-mer exact matches to 472 human genomes (Li, , 2024). 99.7% of them are marked as LC regions by longdust and none of them are alpha satellite. 95.8% of these gorilla-specific regions are distributed within 15Mb from telomeres, broadly in line with Yoo et al., (2025).
TANTAN with “-w500 -r1e-9” reported 672.2Mb of LC regions, 629.1Mb of which overlapped with longdust. TANTAN still tends to report fragmented regions with an N50 block length of 3.7kb. We also ran TRF on the gorilla genome. It did not finish in a week with option “-l20”.
@@ 4 Discussions
Du et al., (2025) observed that each tandem repeat finding tool reported a large fraction of regions unique to itself. We think this is driven by weak tandem repeats that are composed of a few diverged repeat units. If we focus on highly repetitive LC regions, longdust, TRF, TANTAN and ULTRA agree reasonably well with each other. It is difficult to tell which tool is more accurate in terms of genome-wide coverage. We do not intend to brand longdust as the best tool for LC finding; we emphasize more on a new probabilistic model for defining string complexity, which advances the SDUST algorithm and achieves a descent balance between performance, coverage, block contiguity and strand symmetry.
@@
In comparison to tandem repeat finding tools such as TRF and ULTRA, longdust is unable to report the repeat units in case of tandem repeats. This will limit the use cases of longdust. Another practical limitation of longdust is the restricted window size. The genome of Woodhouse’s scrub jays, for example, contains satellite with a 18kb repeat unit (Edwards et al., , 2025). This would be missed by longdust under the default setting. Increasing the window size would make longdust considerably slower. This is partly due to the $O(wL)$ time complexity and partly due to the speedup strategies in longdust that are more effective given $\ell(x)\ll 4^{k}$. It would be ideal to have an algorithm that remains efficient given large windows or, better, does not require specified window sizes.
@@ Acknowledgments
We are grateful to Qian Qin for evaluating the effect of low-complexity regions in structural variant calling, and to Maximilian Haeussler for explaining the role of masking low-complexity regions in cross-species alignment. We thank the reviewers for suggesting cleaner derivation.